# Weyl-Einsteinian-Cubic Gravity

### v1

@@ Abstract
We construct a Weyl-gauging of strongly coupled Einsteinian-Cubic gravity (ECG) in Weyl’s geometry via upgraded Weyl’s curvature tensors, which include abelian gauge fields, and properly tuned compensating real scalar fields. The model is free from any dimensionful parameters. The bare ECG emerges as the lower energy limit of the Weyl-ECG in the local non-conformal-invariant vacua (i.e., broken phase) in the maximally symmetric spacetimes fixing the vacuum expectation value of the scalar field to the Planck mass scale. Here, the natural presence of (anti-) de Sitter backgrounds spontaneously breaks Weyl’s local conformal symmetry akin to the Higgs mechanism, while it is radiatively broken at the renormalization scale at the one-loop level in flat vacua through the Coleman-Weinberg mechanism. The model allows anti-de Sitter and flat spaces but does not allow de Sitter to be vacuum spacetime solutions. The properties of the model deserve further exploration, specifically, those of nonperturbative (e.g., instantons and/or anti-instantons) contributions, for example, in the resurgence or tachyon condensation context requires detailed study.
@@ I Introduction
The biggest obstacle to a UV-complete quantum field theory of general relativity is its inability to achieve unitarity and renormalizability simultaneously. Recall that, during the first and second corrections to Compton scattering due to graviton loops at the radiative scale in general relativity within the context of perturbative quantum field theory, the vertices turn out to be exclusively momenta-dependent, unlike quantum electrodynamics. Consequently, quartic infinities arise at the one-loop level, and sextic infinities emerge at the two-loop level at higher momenta in the UV limit. The renormalization of quantum general relativity unavoidably requires an infinite number of counterterms in the compact form as $\triangle S\sim\sum_{n=0}^{\infty}\int d^{4}x\,(\kappa^{2}R)^{n}R^{2},$ where $\kappa$ is the dimensionful Newton’s constant. As is known, the one-loop divergence is fixed through the topological Gauss-Bonnet term and a particular field redefinition. But the existing catastrophic infinities at the two-loop cannot be renormalized. Further, the theory gets out of control even at the one-loop level as one takes into account all the possible internal matter (scalars, spinors, photons, Yang-Mills) loops, too. Thus, general relativity is non-renormalizable, albeit being unitary [1, 2]. It has been demonstrated in [3] that general relativity integrated by specific quadratic curvature combinations turns out to be power-counting renormalizable despite losing unitarity this time. Having a complete quantum theory of gravity inevitably necessitates the remedy of its very fundamentals [e.g., (loop-level) propagator structure etc.] by assuming either extra appropriate symmetries or higher-order modifications in curvature (with or without other viable lower/higher spin fields) etc., such that those annexed parts would be effective at the sufficiently higher frequencies but suppressed in the lower frequencies. Since the renormalization of general relativity requires infinite numbers of higher-curvature counter terms at each high energy scale, and also since it was shown in [3] that the quadratic curvature modification absorbs the existing divergences despite the cost of losing unitarity, the higher-order curvature gravity models (with or without extra lower spin fields) have occupied a crucial place in this research area. In this respect, as the standout and relevant alternative approaches in the last decade, see, e.g., [4, 5, 6] for how the propagator structure of general relativity gets modified by non-local infinite derivative curvature interactions through particular form factors putting forward intriguing explanations at both the classical and quantum levels. Higher-order curvature terms also arise in string theory and holography, such as specific quadratic and cubic curvature corrections in the low energy limits of distinct types of string theories [7, 8] or quasi-topological higher-derivative curvature corrections in holography [9, 10, 11, 12, 13, 14, 14]). As toy models, higher-curvature modifications have also been widely studied in lower dimensions for a long time. Among those, despite its so-called problems in the holographic perspective at the chiral points, the most successful one is the Topologically Massive Gravity (TMG) by Deser, Jackiw, and Templeton which contains the higher-derivative gravitational Chern-Simons topological term [16, 17].
@@
As is naturally expected, any viable covariant higher curvature gravity model must have the following characteristics: ($a$) The linearized particle content around the constant curvature vacua are identical to those of ordinary general relativity, namely, massless transverse spin-$2$ along with the extra emerging propagating degrees of freedom such as massive spin-$2$ or spin-00 mode. ($b$) The existing relative couplings among the compensating higher-order curvature invariant terms are dimension-independent. ($c$) It must not be trivial or topological in $4D$ for the sake of a dynamical theory. For the last decade, despite its shortcomings such as existing ghost modes (instabilities), a particular cubic curvature gravity model entitled “Einsteinian Cubic Gravity (ECG)” has been proposed in [18]. ECG is interesting since it transcends the so-called Gauss-Bonnet or cubic-Lovelock modifications and fulfills all three prerequisites aforesaid and thus turns out to also be a nontrivial and non-topological cubic-curvature model, unlike the cubic Lovelock model [18, 19]. (See [20, 21, 22, 23, 24, 25, 26, 27, 28] for the static, stationary, and charged modified black hole solutions of ECG and [29, 30] for its quasi-topological generalizations. For the cosmological early and late expansion epochs as well as a smooth universe bouncing between two de Sitter vacua in ECG, see [31, 32, 33, 34, 19]. For the studies on quasinormal modes, gravitational shadow and lensing in ECG, see [35, 36, 37]. See [38] for the Braneworlds in ECG and [39] for a comprehensive thesis in general on geometric modifications of gravity.) But, as is known, since ECG does not belong to the Lovelock family, it contains Ostrogradski ghost-like modes about arbitrary vacua as anticipated and involves pathologies [40, 40, 42]. First of all, one should notice that those instabilities turn out to be catastrophic as one approaches the theory as a UV-complete model, whereas it is secure as to the perturbative perspective via cautious fine-tuning of couplings. In this case, one can study that either the ghost modes are infinitely massive and hence not propagating degrees of freedom such that it decouples in strongly coupled regimes about some particular spacetimes and thus resolves the existing instability [42] [^1] In fact, it is expected to have similar pathologies as extra (massive) degrees of freedom are considered [43, 44]. In such cases, e.g., with the notable Vainshtein radius, [45] shows that nonlinear effects could repress the problematic degrees of freedom in the linearized model at the perceptible scales. Of course, this or other alternative approaches can be addressed to render the ghost modes to proceed toward a UV-complete ECG that is already safe to be studied at the perturbative scale. Due to its notable exclusive attributes mentioned above, ECG surely deserves to be further studied, at least as an effective theory. Furthermore, as for the failures encountered in the full quantum ECG, it is indeed expected to have such problems as one considers ECG as a UV-complete theory since it involves general relativity in its core. That is, the main obstacle against a well-behaved quantum theory of gravity is the dimensionful nature of Newton’s constant (i.e., $[\kappa]=m^{-2}$) which obliges the assumption of infinite numbers of counter-terms at each energy scale for such a desired legitimate complete quantum gravity. Thus, it is natural to anticipate that such a UV-complete theory be free from dimensionful parameters as one moves toward the high-frequency domains, (e.g., Planck’s length) [^2] Note that it is consistent with the eminent “Dirac large numbers hypothesis” stating that the universal constants are not constant at all, rather they rely on the distinct energy scales of the Universe [46, 47] as well as the Hermann Weyl’s proposal in [48, 49, 50].. Actually, this is also the case in special relativity, that is, the particles’ masses do not retain their significance for sufficiently higher energies. Thus, all these suggest boosting Newton’s constant to a viable field augmented with the local conformal symmetry for generic spacetimes.
@@
In addition to all those observations about ECG, recall the following attribute of gravity once more: in general relativity with the torsionless and metric compatibility restrictions, the metric becomes the sole dynamical quantity controlling the sort of geometry and affine structure of the substance manifolds. So, the degrees of freedom associated with the torsion and non-metricity are omitted, and thereby the spacetimes with the Levi-Civita connections are taken as the only physical solutions. However, a broader geometrical description of gravity seems to be essential to shed light on such a desired ultimate theory which will affirmatively upgrade the basic characteristics of general relativity even deeper into extremely high energy regimes (i.e., Planck scale and beyond) where its regular form breaks down [51, 52, 53, 54]. Having a broader geometrical description of the ordinary general relativity without assuming any additional modification will require fully or partially lifting the limitations (namely, torsion-free and metricity) on the connection which were initially set for the sake of metric compatibility. Due to the points mentioned so far and many other reasons in the experimentally confirmed standard model gauge theories, the partial relaxation of metricity restriction in the regular general relativity predominantly deserves more attention as it supplies a natural geometrical environment for the local scale-invariant field theories. This upgrading of gravitational connection is dubbed “Weyl’s conformal gauge modification”. Here, the so-called global rigid scale invariance mandating conformal flatness for Lorentz invariance [^3] For a given field theory involving the specific configuration of field operator $\widehat{\varphi}$ to be scale-invariant (i.e., conformal to a complex plane) under the transformations $x\rightarrow x^{{}^{\prime}}=\lambda x$ only if $\widehat{\varphi}\rightarrow\widehat{\varphi}(x{{}^{\prime}})=\lambda^{h}\widehat{\varphi}(x)$ where $\lambda$ is any constant and $h$ is the scaling dimension of field operator. is boosted to a local scale invariance to achieve the Poincaré invariant theories in any curved spacetimes [48, 49, 50, 55, 56, 57, 58, 59]. In fact, in a parallel analogy, for instance, with the standard model Higgs mechanism, it is affirmative to expect the local scale symmetry to be one of the fundamental symmetries of nature, at least at the UV regimes, because it does not tolerate any dimensionful couplings (say, Newton’s constant, Higgs mass) which can be generated only after the local scale invariance is broken through a reasonable mechanism. See, e.g., [60, 61] for how the local scale symmetry is coherently incorporated into the standard model and [62] for a recent study where it is shown that the metric-affine gravity provides a compatible particle physics. See [63, 64] for the one-loop beta functions in Weyl-invariant gravity models and [65] for how to quantize a Weyl-invariant gravity model coupled to a Stuckelberg photon through background field method. See [66] where Jackiw and Grumiller showed that the Einstein-Weyl equations emerge as a consequence of (anti-)self-duality of $3+1$-dimensional Weyl tensor in the Kaluza-Klein reduction from four to three dimensions. See also [67] where it has been demonstrated that the related asymptotic symmetry generating Lee-Wald charges of Weyl-invariant TMG [68] are both closed and integrable wherein the associated charge algebra is composed of a class of supertranslation, two families of Witt algebras and a family of multiple charges. For a generic enhanced Gauss-Bonnet gravity theory in Weyl’s geometry giving the vector-tensor Horndeski interaction in $4-$dimensions, see [69]. Finally, see, e.g., [70] for the n-dimensional extension of quadratic curvature gravity in Weyl’s geometry, and [71, 72, 73, 74] for some other related works.
@@
Combining all the above-mentioned crucial points and observations, we will formulate a $3+1$-dimensional Weyl-invariant ECG in Weyl’s geometry with the help of appropriately tuned weights of compensating real scalar fields and Weyl’s curvature tensors composed of both the regular Riemannian curvature tensors and abelian gauge fields in this paper. We will see that the model does not involve any dimensionful parameter and the regular ECG comes about in the local non-conformal-invariant vacua (i.e., broken phase) of the model in the maximally symmetric spacetimes. Here, Weyl’s local conformal symmetry will be spontaneously broken akin to the Higgs mechanism in (anti)-de Sitter spaces and radiatively broken at one-loop level in flat spaces via the notable Coleman-Weinberg mechanism. We will also see that the anti-de Sitter and flat spaces are permitted but de Sitter is excluded to be vacuum spacetime solutions.
@@
The layout of the paper is as follows: In Sec. II, we present the preliminaries necessary during calculations. In Sec. III, we formulate the Weyl-Einsteinian-Cubic gravity and introduce its fundamental characteristics. Sec. IV is devoted to the derivation of the field equations for the dynamical fields for arbitrary background spacetimes. In Sec. V, we study the field equations in the maximally symmetric vacuum spacetimes and the corresponding symmetry-breaking mechanism. In Sec. VI, we conclude our results and examine the prospective road maps. Lastly, all the essential expressions required to analyze and understand the model are given in the Appendix.
@@ II Preliminaries
Due to its nature as well as the explicit reasons we present below, it is a priori natural to expect a complicated action describing Weyl-Einsteinian-Cubic gravity (WECG). As we shall see, the explicit version of WECG includes not only the bare cubic curvature terms but also additional higher-order curvature terms up to second order, nonminimally coupled to gauge and scalar fields, as well as self-interacting gauge and scalar field terms up to sextic and quartic order, respectively. For this reason, before dwelling on the details of the model, let us now present some crucial preliminaries that supply convenience during calculations. Firstly, our metric signature convention is (,+,+,+)(-,+,+,+)( - , + , + , + ) and we use $[\nabla_{\mu},\nabla_{\nu}]V_{\alpha}=R_{\mu\nu\alpha\beta}V^{\beta}$. Secondly, let us also define two classes of two-fold indices operations which will be very useful particularly during the derivations of the field equations for $(\delta g^{\mu\nu},\delta A^{\mu},\delta\Phi)$: the first one is called as the “two-fold symmetric-(anti)symmetric indices operations” providing the following operations between indices
@@
III Weyl-Einsteinian Cubic Gravity

### v2

# Cosmological Weyl-Einsteinian-Cubic Gravity as a Gauge Theory of Gravity

@@ Abstract
We construct a Weyl-Einsteinian-Cubic Gravity (ECG) as a cubic gauge theory of gravity via abelian gauge and properly tuned compensating real scalar fields. The model is free from any dimensionful parameters. The bare ECG emerges as the lower energy limit of the Weyl-ECG in the local non-conformal-invariant vacua (i.e., broken phase) in the maximally symmetric spacetimes fixing the vacuum expectation value of the scalar field to the Planck mass scale. Here, the natural presence of (anti-) de Sitter backgrounds spontaneously breaks Weyl’s local conformal symmetry akin to the Higgs mechanism, while it is radiatively broken at the renormalization scale at the one-loop level in flat vacua through the Coleman-Weinberg mechanism. The model allows anti-de Sitter and flat spaces but does not allow de Sitter to be vacuum spacetime solutions. The properties of the model deserve further exploration, specifically, those of nonperturbative (e.g., instantons and/or anti-instantons) contributions, for example, in the resurgence or tachyon condensation context requires detailed study.
@@ I Introduction
The biggest obstacle to a UV-complete quantum field theory of general relativity is its inability to achieve unitarity and renormalizability simultaneously. Recall that, during the first and second corrections to Compton scattering due to graviton loops at the radiative scale in general relativity within the context of perturbative quantum field theory, the vertices turn out to be exclusively momenta-dependent, unlike quantum electrodynamics. Consequently, quartic infinities arise at the one-loop level, and sextic infinities emerge at the two-loop level at higher momenta in the UV limit. The renormalization of quantum general relativity unavoidably requires an infinite number of counter-terms in the compact form as $\triangle S\sim\sum_{n=0}^{\infty}\int d^{4}x\,(\kappa^{2}R)^{n}R^{2},$ where $\kappa$ is the dimensionful Newton’s constant. As is known, the one-loop divergence is fixed through the topological Gauss-Bonnet term and a particular field redefinition. But the existing catastrophic infinities at the two-loop cannot be renormalized. Further, the theory gets out of control even at the one-loop level as one takes into account all the possible internal matter (scalars, spinors, photons, Yang-Mills) loops, too. Thus, general relativity is non-renormalizable, albeit being unitary [1, 2]. It has been demonstrated in [3] that general relativity integrated by specific quadratic curvature combinations turns out to be power-counting renormalizable despite losing unitarity this time. Having a complete quantum theory of gravity inevitably necessitates the remedy of its very fundamentals [e.g., (loop-level) propagator structure etc.] by assuming either extra appropriate symmetries or higher-order modifications in curvature (with or without other viable lower/higher spin fields) etc., such that those annexed parts would be effective at the sufficiently higher frequencies but suppressed in the lower frequencies. Since the renormalization of general relativity requires infinite numbers of higher-curvature counter terms at each high energy scale, and also since it was shown in [3] that the quadratic curvature modification absorbs the existing divergences despite the cost of losing unitarity, the higher-order curvature gravity models (with or without extra lower spin fields) have occupied a crucial place in this research area. In this respect, as the standout and relevant alternative approaches in the last decade, see, e.g., [4, 5, 6] for how the propagator structure of general relativity gets modified by non-local infinite derivative curvature interactions through particular form factors putting forward intriguing explanations at both the classical and quantum levels. Higher-order curvature terms also arise in string theory and holography, such as specific quadratic and cubic curvature corrections in the low energy limits of distinct types of string theories [7, 8] or quasi-topological higher-derivative curvature corrections in holography [9, 10, 11, 12, 13, 14, 15]). As toy models, higher-curvature modifications have also been widely studied in lower dimensions for a long time. Among those, despite its so-called problems in the holographic perspective at the chiral points, the most successful one is the Topologically Massive Gravity (TMG) by Deser, Jackiw, and Templeton which contains the higher-derivative gravitational Chern-Simons topological term [16, 17].
@@
As is naturally expected, any viable covariant higher curvature gravity model must have the following characteristics: ($a$) The linearized particle content around the constant curvature vacua are identical to those of ordinary general relativity, namely, massless transverse spin-$2$ along with the extra emerging propagating degrees of freedom such as massive spin-$2$ or spin-0 mode. ($b$) The existing relative couplings among the compensating higher-order curvature invariant terms are dimension-independent. ($c$) It must not be trivial or topological in $4D$ for the sake of a dynamical theory. For the last decade, despite its shortcomings such as existing ghost modes (instabilities), a particular cubic curvature gravity model entitled “Einsteinian Cubic Gravity (ECG)” has been proposed in [18]. ECG is interesting since it transcends the so-called Gauss-Bonnet or cubic-Lovelock modifications and fulfills all three prerequisites aforesaid and thus turns out to also be a nontrivial and non-topological cubic-curvature model, unlike the cubic Lovelock model [18, 19, 20]. (See [21, 22, 23, 24, 25, 26, 27, 28, 29] for the static, stationary, and charged modified black hole solutions of ECG and [30, 31] for its quasi-topological generalizations. For the cosmological early and late expansion epochs as well as a smooth universe bouncing between two de Sitter vacua in ECG, see [32, 33, 34, 35, 20]. For the studies on quasinormal modes, gravitational shadow and lensing in ECG, see [36, 37, 38]. See [39] for the Braneworlds in ECG and [40] for a comprehensive thesis in general on geometric modifications of gravity.) But, as is known, since ECG does not belong to the Lovelock family, it contains Ostrogradski ghost-like modes about arbitrary vacua as anticipated and involves pathologies [41, 42, 43]. First of all, one should notice that those instabilities turn out to be catastrophic as one approaches the theory as a UV-complete model, whereas it is secure as to the perturbative perspective via cautious fine-tuning of couplings. In this case, one can study that either the ghost modes are infinitely massive and hence not propagating degrees of freedom such that it decouples in strongly coupled regimes about some particular spacetimes and thus resolves the existing instability [43] [^1] In fact, it is expected to have similar pathologies as extra (massive) degrees of freedom are considered [44, 45]. In such cases, e.g., with the notable Vainshtein radius, [46] shows that nonlinear effects could repress the problematic degrees of freedom in the linearized model at the perceptible scales. Of course, this or other alternative approaches can be addressed to render the ghost modes to proceed toward a UV-complete ECG that is already safe to be studied at the perturbative scale. Due to its notable exclusive attributes mentioned above, ECG surely deserves to be further studied, at least as an effective theory. Furthermore, as for the failures encountered in the full quantum ECG, it is indeed expected to have such problems as one considers ECG as a UV-complete theory since it involves general relativity in its core. That is, the main obstacle against a well-behaved quantum theory of gravity is the dimensionful nature of Newton’s constant (i.e., $\kappa]=m^{-2}$) which obliges the assumption of infinite numbers of counter-terms at each energy scale for such a desired legitimate complete quantum gravity. Thus, it is natural to anticipate that such a UV-complete theory be free from dimensionful parameters as one moves toward the high-frequency domains, (e.g., Planck’s length) [^2] Note that it is consistent with the eminent “Dirac large numbers hypothesis” stating that the universal constants are not constant at all, rather they rely on the distinct energy scales of the Universe [47, 48] as well as the Hermann Weyl’s proposal in [49, 50, 51].. Actually, this is also the case in special relativity, that is, the particles’ masses do not retain their significance for sufficiently higher energies. Thus, all these suggest boosting Newton’s constant to a viable field augmented with the local conformal symmetry for generic spacetimes.
@@
In addition to all those observations about ECG, recall the following attribute of gravity once more: in general relativity with the torsionless and metric compatibility restrictions, the metric becomes the sole dynamical quantity controlling the sort of geometry and affine structure of the substance manifolds. So, the degrees of freedom associated with the torsion and non-metricity are omitted, and thereby the spacetimes with the Levi-Civita connections are taken as the only physical solutions. However, a broader geometrical description of gravity seems to be essential to shed light on such a desired ultimate theory which will affirmatively upgrade the basic characteristics of general relativity even deeper into extremely high energy regimes (i.e., Planck scale and beyond) where its regular form breaks down [52, 53, 54, 55]. Having a broader geometrical description of the ordinary general relativity without assuming any additional modification will require fully or partially lifting the limitations (namely, torsion-free and metricity) on the connection which were initially set for the sake of metric compatibility. Due to the points mentioned so far and many other reasons in the experimentally confirmed standard model gauge theories, the partial relaxation of metricity restriction in the regular general relativity predominantly deserves more attention as it supplies a natural geometrical environment for the local scale-invariant field theories. This upgrading of gravitational connection is dubbed “Weyl’s conformal gauge modification”. Here, the so-called global rigid scale invariance mandating conformal flatness for Lorentz invariance [^3] For a given field theory involving the specific configuration of field operator $\widehat{\varphi}$ to be scale-invariant (i.e., conformal to a complex plane) under the transformations $x\rightarrow x^{{}^{\prime}}=\lambda x$ only if $\widehat{\varphi}\rightarrow\widehat{\varphi}(x{{}^{\prime}})=\lambda^{h}\widehat{\varphi}(x)$ where $\lambda$ is any constant and $h$ is the scaling dimension of field operator. is boosted to a local scale invariance to achieve the Poincaré invariant theories in any curved spacetimes [49, 50, 51, 56, 57, 58, 59, 60]. Notice also that Weyl’s gauging provides not only local conformal symmetry but also a gauge theory of gravity as in the Standard Model of particles rooted on the internal gauge group symmetries444 In fact, in a parallel analogy, for instance, with the standard model Higgs mechanism, it is affirmative to expect the local scale symmetry to be one of the fundamental symmetries of nature, at least at the UV regimes, because it does not tolerate any dimensionful couplings (say, Newton’s constant, Higgs mass) which can be generated only after the local scale invariance is broken through a reasonable mechanism.. See, e.g., [61, 62] for how the local scale symmetry is coherently incorporated into the standard model and [63] for a recent study where it is shown that the metric-affine gravity provides a compatible particle physics. See [64, 65] for the one-loop beta functions in Weyl-invariant gravity models and [66] for how to quantize a Weyl-invariant gravity model coupled to a Stuckelberg photon through background field method. See [67] where Jackiw and Grumiller showed that the Einstein-Weyl equations emerge as a consequence of (anti-)self-duality of $3+1$-dimensional Weyl tensor in the Kaluza-Klein reduction from four to three dimensions. See also [68] where it has been demonstrated that the related asymptotic symmetry generating Lee-Wald charges of Weyl-invariant TMG [69] are both closed and integrable wherein the associated charge algebra is composed of a class of supertranslation, two families of Witt algebras and a family of multiple charges. For a generic enhanced Gauss-Bonnet gravity theory in Weyl’s geometry giving the vector-tensor Horndeski interaction in 44-dimensions, see [70]. Finally, see, e.g., [71] for the n-dimensional extension of quadratic curvature gravity in Weyl’s geometry, and [72, 73, 74, 75, 76] for some other related works.
As is mentioned above, observe that the Standard Model of particles is predominantly established as a gauge theory augmented with the internal compact Lie group symmetries [77, 78]. This has ended up with several experimentally revealed inventions, including the exceptional Higgs mechanism for the generation of the existing masses of the particles through a legitimate symmetry-breaking mechanism. Hence, almost all the fundamental interactions in universe seem to be underlined by the gauge field paradigm, whereas gravity stands out as a curious and notable exclusion. Corollary to this, a reasonable gauge theory of gravity seems to be essential. See particularly [79] (and references therein) for a recent and a comprehensive study on the physical reasons why the upgrading of the paradigm is required, especially why nonmetricity is essential to further shed light on high energy regimes beyond Standard Model.
@@
Combining all the above-mentioned crucial points and observations, we will formulate a $3+1$-dimensional Weyl-invariant ECG in Weyl’s geometry-as a cubic gauge theory of gravity-with the help of appropriately tuned weights of compensating real scalar fields and Weyl’s curvature tensors composed of both the regular Riemannian curvature tensors and abelian gauge fields in this paper. We will see that the model does not involve any dimensionful parameter and the regular ECG comes about in the local non-conformal-invariant vacua (i.e., broken phase) of the model in the maximally symmetric spacetimes. Here, Weyl’s local conformal symmetry will be spontaneously broken akin to the Higgs mechanism in (anti)-de Sitter spaces and radiatively broken at one-loop level in flat spaces via the notable Coleman-Weinberg mechanism. We will also see that the anti-de Sitter and flat spaces are permitted but de Sitter is excluded to be vacuum spacetime solutions.
@@
The layout of the paper is as follows: In Sec. II, we present the preliminaries necessary during calculations. In Sec. III.2, we formulate the Weyl-Einsteinian-Cubic gravity and introduce its fundamental characteristics. Sec. IV is devoted to the derivation of the field equations for the dynamical fields for arbitrary background spacetimes. In Sec. V, we study the field equations in the maximally symmetric vacuum spacetimes and the corresponding symmetry-breaking mechanism. In Sec. VI, we conclude our results and examine the prospective road maps. Lastly, all the essential expressions required to analyze and understand the model are given in the Appendix.
@@ II Preliminaries
Before progressing further, as is explicitly shown in the rest of paper, it is necessary to emphasize here that the Weyl-Einsteinian-Cubic gravity (WECG) provides not only the Poincaré invariant ECG in any arbitrary curved spacetimes but also how the ordinary ECG emerges as the vacuum (lower-energy) limit of WECG as a result of a viable symmetry breaking mechanism (namely, either spontaneously in the (Anti)-de Sitter spacetimes like the Standard Model Higgs Mechanism or radiatively at one-loop level in flat backgrounds as in the Coleman-Weinberg mechanism for the quantum electrodynamics of a scalar field in four-dimensions). One should also note that since the Weyl’s gauge symmetry is achieved with the help of extra gauge and scalar fields, extra tachyonic modes may emerge. But, as it is clearly explained in the conclusion, analogous phenomena (i.e., unstable tachyonic modes with imaginary masses and so vacuum expectation values) also emerge in $D\bar{D}$-brane pairs in string field theory, ending up with the rolling down of the field to a stable vacuum ground state that are free from any observable tachyonic mode; or, in the resurgence theory which combines the perturbative and nonperturbative contributions such that the two-fold ambiguous imaginary part of the Borel resummation of the perturbation theory and the relevant part of instanton-anti-instanton amplitude completely cancel each others leading to a stable system. (For a related detailed discussion and also possible future direction, see the conclusion.) One shall also note that, we will study the foremost fundamentals of model (e.g., construction of the action, field equations, vacuum analysis, etc.) in this work, but of course, since the model suggests intriguing outcomes and all the analyses cannot be done in a single paper, further extra studies are apparently required to understand WECG in all essential aspects.
As for the quantitative base of the model, let us first note that, as in the explicit reasons given below, it is a priori natural to expect a highly complicated action describing WECG since, as is given in (10), its explicit form contains various non-minimal interactions among higher-order curvatures, vector and scalar fields as well as their higher-order self-interactions. As we shall see, the explicit version of WECG includes not only the bare cubic curvature terms but also additional higher-order curvature terms up to second order, nonminimally coupled to gauge and scalar fields, as well as self-interacting gauge and scalar field terms up to sextic and quartic order, respectively. For this reason, before dwelling on the details of the model, let us now present some crucial preliminaries that supply convenience during calculations. Firstly, our metric signature convention is (,+,+,+)(-,+,+,+) and we use $\nabla_{\mu},\nabla_{\nu}]V_{\alpha}=R_{\mu\nu\alpha\beta}V^{\beta}$. Secondly, let us also define two classes of two-fold indices operations which will be very useful particularly during the derivations of the field equations for $(g^{\mu\nu},A^{\mu},\Phi)$: the first one is called as the “two-fold symmetric-(anti)symmetric indices operations” providing the following operations between indices
@@
III Weyl-Einsteinian Cubic Gravity: Construction
III.1 Cosmological Einsteinian Cubic Gravity: A Review
Due to being the base model, let us shortly recall that the usual ECG is promoted to arbitrary cosmological scenarios (FLRW) with additional cubic terms in the Cosmological Einsteinian Cubic Gravity (Cosmological ECG) [19]. While keeping the intriguing characteristics of usual ECG intact, the Cosmological ECG also provides a well-defined initial value problem and further a geometric mechanism driving an inflationary period in the radiation dominated region to a late time cosmology converging to $\Lambda CDM$. In doing so, they start with the plausible cubic combinations
$\begin{array}[]{ll}\mathcal{L}_{1}={{{R_{a}}^{c}}_{b}}^{d}{{{R_{c}}^{e}}_{d}}^{f}{{{R_{e}}^{a}}_{f}}^{b}\leavevmode\nobreak\ ,&\mathcal{L}_{2}={R_{ab}}^{cd}{R_{cd}}^{ef}{R_{ef}}^{ab}\leavevmode\nobreak\ ,\\[4.29993pt] \mathcal{L}_{3}=R_{abcd}{R^{abc}}_{e}R^{de}\leavevmode\nobreak\ ,&\mathcal{L}_{4}=R_{abcd}R^{abcd}R\leavevmode\nobreak\ ,\\[4.29993pt] \mathcal{L}_{5}=R_{abcd}R^{ac}R^{bd}\leavevmode\nobreak\ ,&\mathcal{L}_{6}={R_{a}}^{b}{R_{b}}^{c}{R_{c}}^{a}\leavevmode\nobreak\ ,\\[4.29993pt] \mathcal{L}_{7}=R_{ab}R^{ab}R\leavevmode\nobreak\ ,&\mathcal{L}_{8}=R^{3}\leavevmode\nobreak\ ,\\[4.29993pt] \mathcal{L}_{9}=\nabla_{a}R_{bc}\nabla^{a}R^{bc}\leavevmode\nobreak\ ,&\mathcal{L}_{10}=\nabla_{a}R\,\nabla^{a}R\leavevmode\nobreak\ .\end{array}$ (3)
Referring [19] for details, let us quote their results of that due to the topological effects and the requirement that it shall preserve all the promising characteristics of ECG, one ends up with the following particular cubic combinations
$\mathcal{P}=12\mathcal{L}_{1}+\mathcal{L}_{2}-12\mathcal{L}_{5}+8\mathcal{L}_{6}\leavevmode\nobreak\ ,\qquad\mathcal{C}=\mathcal{L}_{3}-\frac{1}{4}\mathcal{L}_{4}-2\mathcal{L}_{5}+\frac{1}{2}\mathcal{L}_{7}\leavevmode\nobreak\ ,$ (4)
which describes a unitary massless graviton around maximally symmetric vacua. Observe that the terms $\mathcal{L}_{8}=R^{3},\mathcal{L}_{9}=\nabla_{a}R_{bc}\nabla^{a}R^{bc}$ and $\mathcal{L}_{10}=\nabla_{a}R\,\nabla^{a}R$ in 3 drop in the construction of the Cosmological ECG. Moreover, it has been shown in [19] that, in addition to preserving all the promising characteristics of ECG, a particular combination of $\mathcal{P}$ and $\mathcal{C}$ also provides a cosmological scenario (in FLRW ansatz) wherein the set of Friedmann equations remains second order.
III.2 Weyl-Invariant Extension of Einsteinian Cubic Gravity
Now that we have briefly presented the preliminaries for our construction, let us now proceed with our focal purpose of concretizing the Weyl-invariant extention of the Cosmological ECG which we abbreviate as WECG. First of all, note that we construct the Weyl’s gauging of the particular cubic model of the Cosmological ECG [19, 20] and thus take the specific cubic curvature terms in (4). Accordingly, it will be primarily expected from the WECG to recover the main attributes of the ordinary Cosmological ECG in the subregion of its broken phase. At this point, one shall observe that the main projective of the paper is to build the Weyl-invariant extension of the above-mentioned particular cubic model, namely, the Cosmological ECG which describes $7$ terms and so $7$ generic couplings. But, of course, this is a very active research fields and hence there are other alternative directions to shed further lights on the local conformal invariant cubic curvature theories. Here, having a viable Weyl-invariant extension of the other cubic models based on the Weyl tensors and traceless Ricci tensor is particularly interesting because it has potential to provide less number of independent higher-derivative terms and so relatively shorter field equations: in this regard, e.g., see [80] which posits that there are $6$ independent terms in $3+1$ dimensions or [81] (and more explicitly in (3.6) of [82]) for a $6$ dimensional conformal quantum gravity describing certain cubic terms of Weyl tensors, Ricci tensors and particular divergence terms. At this point, see also [83, 84] for another interesting alternative conformal quantum gravity model constructed with the help of (nonlocal) higher derivative gravitational corrections wherein intriguing outcomes such as resolutions of catastrophic problems towards a complete theory are presented. Observe that the Weyl’s gauge invariant cubic gravity that we shall construct and study, in general, is different in many aspects: in our construction, we shall take the basics (particularly, the local gauge field nature) of the $3+1$-dimensional Standard model of fundamental particles as one of guiding (reference) paths. This is mainly because of the following main reasons: firstly, as was mentioned in the introduction, the Standard model has been verified by countless number of experiments and almost all the fundamental interactions are based on the gauge field paradigm, while gravity stands out as a curious and notable exclusion. So, a reasonable gauge theory of gravity seems to be essential. (See, e.g., [79] [and references therein] for a recent intriguing study on the requirement of such a gauge theory of gravity.) [^5] Observe that despite many appealing higher dimensional models, since the observed universe is $3+1$, we mainly focus on the construction of a complete well-behaved $3+1$ dimensional gauge gravity model.. Thus, as in the Standard model gauge theory, the WECG that we dwell on is achieved with the help of extra abelian gauge field unlike those in [81, 82, 83, 84]. Hence, this construction supplies a gauge theory of gravity model analogous to Standard model gauge theories. This approach of local conformal-invariance also provides Poincare invariant theories for not only certain curved spacetimes having asymptotical conformal group but also any arbitrary spacetimes, whereas those approaches, e.g., in [81, 82, 83, 84] do not assume any extra gauge field and they require the spacetime solutions to be asymptotically conformal invariant [^6] For example, note that the construction in [81] requires asymptotically (A)dS spacetime solution together with the selection of some certain Neumann boundary conditions for the sake of viable limits. Of course, since all those distinct approaches are notable ones and deserves to be studied in all possible aspects, we suggest all of them to get further possible insights towards a ultimate theory.